Write each quadratic function in the form f(x) = a (x-h)^2 + k
F(x)= -3x^2 - 12x - 13

Found 2 solutions by nerdybill, DrBeeee:
Answer by nerdybill(7384)

Write each quadratic function in the form f(x) = a (x-h)^2 + k
F(x)= -3x^2 - 12x - 13
Complete the square:
F(x)= (-3x^2 - 12x) - 13
F(x)= -3(x^2 + 4x) - 13
F(x)= -3(x^2 + 4x + 4) - 13 + 12
F(x)= -3(x+2)(x+2) - 1
F(x)= -3(x+2)^2 - 1

Answer by DrBeeee(684)

Puting F(x) into the form of f(x), is called completing the square.
To do this, start with the first two terms of F(x),
(1) -3x^2 - 12x and factor out the coefficient of x^2 as follows,
(2) -3*{x^2 + 4*x}
This can be shown to equate to
(3) -3*{(x+2)^2 - 4}
We can show this by FOILing (x+2)^2 in (3), and get
(4) -3*{x^2 + 4*x + 4 - 4} or
(5) -3*{x^2 + 4*x}, which is the same as (2)
Therefore we can replace the first two terms of F(x) with (3), and get
(6) F(x) = -3*{(x+2)^2 - 4} - 13, which is exactly the same as the given F(x).
Notice that we have created a shifted variable (x+2)^2 to replace the plain x^2 of F(x) as required by the form of f(x), thus it is called completing the square.
Now simplify (6) and get
(7) F(x) = -3*(x+2)^2 -3*(-4) - 13 or
(8) F(x) = -3*(x+2)^2 - 1
Note that (8) matches the form given by
(9) f(x) = a*(x-h)^2 + k
where
a = -3,
h = -2, and
k = -1
Comment: The neat part about this form of the quadratic is that the point (h,k) is the turning point of the parabola. And x = h is the axis of symmetry!